OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (2, -6) ; Slope: -3/4

OpenStudy (anonymous):

i think the equation you need to plug it into is y=mx+b

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ -6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{3}{4} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

hmmm actually if you need it in "point-slope form" then just don't solve for "y" :) leave it like so, notice that's the point-slope form already anyhow

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I think..

OpenStudy (jdoe0001):

plug in your values... they're all given anyhow

OpenStudy (anonymous):

okay, the first part is y-6 , right?

OpenStudy (anonymous):

I just want to make sure im doing it right

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ -6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{3}{4} \\ \quad \\ y-{\color{blue}{ (-6)}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\\ \qquad \uparrow\\ \textit{point-slope form}\) -(-6) = +6

OpenStudy (anonymous):

y + 6 = - 3/4 (x-2) ?

OpenStudy (jdoe0001):

yeap